The basal edge is 6 cm. of the regular triangular prism. What is the area of the base?


1
Expert's answer
2021-02-23T15:32:52-0500

If the prism is regular then its base is an equilateral triangle with a side of 6 cm. Its area is "\frac{a^2\sqrt3}{4}" , where a is the length of a side. So the area is "\frac{6^2\cdot\sqrt3}{4}=9\sqrt3" "cm^2" .